OpenStudy (anonymous):

need help (x^3+y^2)(x^3-y^2) can someone help me figure this out and explain?

OpenStudy (anonymous):

you have to do four multiplications

OpenStudy (anonymous):

I need to multiply

OpenStudy (anonymous):

\[x^3\times x^3\] is the first one

OpenStudy (anonymous):

x^6? do I add the exponents?

OpenStudy (anonymous):

then \(x^3\times (-y^2)\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then \(x^3\times y^2\)

OpenStudy (anonymous):

and finally \(y^2\times (-y)^2\)

OpenStudy (anonymous):

x^6-y^4?

OpenStudy (anonymous):

you will notice that middle terms add up to zero, so you are only left with \[x^6-y^4\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it is always true that \((a+b)(a-b)=a^2-b^2\)

OpenStudy (anonymous):

see, all I need is confidence and your great guidings! :) I might make it after all. LOL

jimthompson5910 (jim_thompson5910):

you can also use a table to help you multiply